NONEXISTENCE OF CONTINUOUS PEAKING FUNCTIONS

Jiye Yu

Abstract. We construct a smoothly bounded pseudoconvex domain such that every boundary point has a p.s.h. peak function but some boundary point admits no (local) holomorphic peak function.

Typeset by AMS-TEX